OpenStudy (anonymous):

What are the solutions of the system? y = x^2 + 4x - 5 y = 2x - 2

OpenStudy (anonymous):

y=2x-2 so 2x-2=x^2+4x-5

OpenStudy (anonymous):

Get it equal to 0 and factor x^2+2x-3=0

OpenStudy (anonymous):

I solved it by graphing and got -3, -8 and 1, 0

OpenStudy (anonymous):

There should only be -3 and 1

OpenStudy (anonymous):

It doesn't have that choice as an answer. (It's multiple choice)

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

A. (–3, –8) and (1, 0) B. (–3, –8) and (–1, –4) C. (–3, 0) and (1, –8) D. no solution

OpenStudy (anonymous):

I got A by graphing it. So those are the X and Y coordinates

OpenStudy (anonymous):

Oh i didnt realize you were typing coordinates. A is right then

OpenStudy (anonymous):

Thanks for helping me Caleb. ^.^